OpenStudy (anonymous):

find the approximate integration for the interval 3cos(x^2)dx from 0 to 1 when n=4.Use the trapezoid rule

OpenStudy (anonymous):

f(x)=cos(x^2) deltax=1-0/4=1/4=.25 in degrees f(0)=1,f(.25)=.999,f(.5)=.999,f(3/4)=.999,f(1)=.998 so .25/2(1+.999+.999+.999+.999+.999+.999+.9984)=.99905